From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat May 03 2003 - 23:56:21 EDT
------------------------------------------------------------
Date: Sat May  3 23:56:21 2003, Shift: 16:00-24:00
Supervisor: rk, with: 
Log book#14 pages 92 - 92
Shift Summary:
------------------------------------------------------------
1. We had beam almost all shift (two nice stores, which were abruptly stopped by dumping votes of our two Big Brothers)
2. Runs   Tr6  Tr3  other
   8784   31k  30k
   8785   87k  85k
   8786  101k  98k
                    -> FS rotated to 12 deg
   8791   10k  56k  FS @ 12deg @ 1/4B
                    -> FS changed magnets to 1/15B
   8792   30k  69k  FS @ 12deg @ 1/15B
Goal: to collect ~300k with FS @ 8deg @ 1/12B was achieved
Run 8791 is for 'Rate checkout'
3. No new detector problems (TOF II is still not working perfectly)
------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Sat May 03 2003 - 23:56:53 EDT